# ExeCoder: Empowering Large Language Models with Executability Representation for Code Translation

### v2

@@ 4. Experimental Evaluation
4.1. Experimental Setup
Dataset. We conducted comprehensive experiments on the widely used code translation public dataset TransCoder-test (Roziere et al., 2020) to demonstrate the effectiveness of the ExeCoder. TransCoder-test contains solutions to programming problems collected from GeeksForGeeks, with each problem implemented in C++, Java, and Python, totaling 948 parallel samples. TransCoder-test is equipped with predefined unit test templates for over half of the translation pairs to evaluate whether the generated functions return the same output as the reference functions given the same input. Each unit test consists of ten test cases.
However, these unit test templates establish fixed parameter passing methods or return types that can only evaluate specific implementations. When the generated functions do not conform to the predefined implementations, the unit tests cannot execute as expected, even if the generated functions are functionally equivalent to the reference functions, resulting in unreliable evaluation. To address this, we enhanced TransCoder-test to ensure that the unit test results are capable of assessing the code translation ability. The enhanced test set is referred to as TransCoder-test-X, with detailed processing cases provided in Appendix D.
Match-based Metrics. The matching-based metrics aim to evaluate the quality of generated code through static code analysis. Following existing work (Huang et al., 2023), we employed three metrics: Exact Match (EM), BLEU (Papineni et al., 2002), and CodeBLEU (Ren et al., 2020) for evaluation. These metrics can measure the n-gram overlap, syntactic structure, and semantic equivalence between the generated code and the target code.
Execution-based Metrics. The execution-based metrics aim to assess the quality of generated code through the execution status of the code. Following existing work (Roziere et al., 2020), we utilize test computational accuracy (TCA) as a metric, which measures the proportion of code that passes unit tests. Furthermore, to further reflect the fine-grained execution status of the code, we have designed two additional metrics: Compilation Accuracy (CA) and Case Computational Accuracy (CCA), the former measuring the proportion of code that successfully compiles, and the latter measuring the proportion of code that passes case tests.
Baselines. To assess the effectiveness of ExeCoder, we compare it with state-of-the-art LLMs. Specifically, we use three advanced closed-source models as baselines: OpenAI’s GPT-3.5, GPT-4, and GPT-4-o (Achiam et al., 2023). Additionally, we select five leading open-source code LLMs as baselines: Deepseek-Coder-6.7b-instruct (Guo et al., 2024), CodeLLama-7B (Roziere et al., 2023), Magicoder-S-DS-6.7B (Wei et al., 2024), Qwen2.5-Coder-7B (Hui et al., 2024), and Wavecoder-ultra-6.7b (Yu et al., 2024).
Implementation Details. To ensure the reproducibility of the research, we present detailed implementation specifics. Deepseek-Coder-6.7b-instruct serves as our basic foundation model. To fine-tune the base model, we employed specific configurations, including a batch size of 128, a sequence length of 3076, a learning rate of 2e-6, a cosine learning rate scheduler, and fp16 mixed precision.
@@ 4.1. Experimental Setup
Table 2. Match-based evaluation of ExeCoder and baseline models on TransCoder-test-X.
@@
4.2. Experimental Results
Effectiveness of ExeCoder. To assess the effectiveness of ExeCoder in code translation task, we evaluated ExeCoder on the TransCoder-test-X, which includes translations among the three most commonly used programming languages: C++, Python, and Java. Tables 1 and 2 present the comparison results with baseline models.
  • Code: Fine-tuning using only the source code.
Tables 3, 4, 5 and Figure 4 present our results, which indicate that the learning from each executability representation contributes to varying degrees of improvement in the model’s code translation capability. For execution-based metrics, the inclusion of the three different types of executability representations resulted in an average increase of 0.41%, 1.64%, and 1.21% across the three metrics. For match-based metrics, the integration of the three different types of executability representations led to an average increase of 0.39%, 5.56%, and 5.24% across the three metrics. These advantages indicate that executability representations can provide execution status beyond code semantics, enhancing the cross-language understanding capability of LLMs.
@@ 4.2. Experimental Results
Figure 4. Match-based evaluation of different variants of ExeCoder on TransCoder-test-X.
Table 3. Compilation Accuracy (CA) of different variants of ExeCoder.
Model Code with FS with SS with VD
From C++ 91.13 91.56 (+0.43) 91.66 (+0.53) 91.14 (+0.01)
To C++ 96.36 96.47 (+0.11) 96.90 (+0.54) 97.11 (+0.75)
From Python 89.21 89.52 (+0.31) 91.20 (+1.99) 90.78 (+1.57)
To Python 90.41 91.59 (+1.18) 93.00 (+2.59) 92.89 (+2.48)
From Java 92.68 93.54 (+0.86) 95.05 (+2.37) 95.27 (+2.59)
To Java 86.25 86.56 (+0.31) 88.02 (+1.77) 87.19 (+0.94)
Average 91.01 91.54 (+0.53) 92.64 (+1.63) 92.39 (+1.38)
Table 4. Case Computational Accuracy (CCA) of different variants of ExeCoder.
Model Code with FS with SS with VD
From C++ 86.04 86.48 (+0.44) 87.33 (+1.29) 86.56 (+0.52)
To C++ 91.21 90.89 (-0.32) 91.68 (+0.47) 91.78 (+0.57)
From Python 84.27 84.28 (+0.01) 85.37 (+1.10) 84.79 (+0.52)
To Python 83.11 83.96 (+0.85) 85.83 (+2.72) 85.59 (+2.48)
From Java 87.13 87.43 (+0.30) 89.38 (+2.25) 89.38 (+2.25)
To Java 83.11 83.33 (+0.22) 84.57 (+1.46) 83.36 (+0.25)
Average 85.81 86.06 (+0.25) 87.36 (+1.55) 86.91 (+1.10)
Table 5. Test Computational Accuracy (TCA) of different variants of ExeCoder.
Model Code with FS with SS with VD
From C++ 81.62 82.17 (+0.55) 82.90 (+1.28) 82.16 (+0.54)
To C++ 85.22 85.12 (-0.10) 86.19 (+0.97) 86.08 (+0.86)
From Python 78.28 78.60 (+0.32) 80.08 (+1.80) 78.93 (+0.65)
To Python 76.40 77.59 (+1.19) 79.31 (+2.91) 79.20 (+2.80)
From Java 81.82 82.25 (+0.43) 83.98 (+2.16) 84.09 (+2.27)
To Java 80.10 80.31 (+0.21) 81.46 (+1.36) 79.90 (-0.20)
Average 80.58 81.01 (+0.43) 82.32 (+1.74) 81.73 (+1.15)
Effectiveness of Progressive Executability Representation Learning. We conduct an ablation study to evaluate the effectiveness of the PERL strategy. Specifically, we assess the impact of each fine-tuning phase in PERL on the model’s code translation capability. Each fine-tuning phase focuses on learning a specific executability representation, and the next fine-tuning phase begins only after the current one has converged. Tables 6, 7, and 8 present our results.
Overall, ExeCoder’s meticulously designed PERL strategy significantly enhances the code translation performance of LLMs through the learning of executability representations. Compared to models fine-tuned solely on source code, the inclusion of executability representations raises the upper limits of model capability, with three execution-based metrics improving by 0.53%, 0.85%, and 1.28%, respectively. Furthermore, the learning of specific executability representations at each stage improves the model’s cross-language understanding. Relative to the previous stage, each fine-tuning stage further improved the key execution-based metric TCA by 0.50%, 0.30%, and 0.49%, respectively. This evidence indicates that the PERL strategy indeed facilitates a better understanding of executability representation.
Table 6. Compilation Accuracy (CA) of various stages within PERL.
Model Code Code-FS Code-FS-$S$ Code-FS-$S$-VD
From C++ 91.55 91.13 (-0.42) 91.15 (+0.02) 91.56 (+0.41)
To C++ 96.36 96.36 (+0.00) 97.11 (+0.75) 97.00 (-0.11)
From Python 89.73 90.78 (+1.05) 90.58 (-0.20) 91.00 (+0.42)
To Python 93.00 92.56 (-0.44) 93.53 (+0.97) 93.53 (+0.00)
From Java 95.16 94.62 (-0.54) 95.37 (+0.75) 95.48 (+0.11)
To Java 87.08 87.60 (+0.52) 86.46 (-1.14) 87.50 (+1.04)
Average 92.15 92.18 (+0.03) 92.37 (+0.19) 92.68 (+0.31)
Table 7. Case Computational Accuracy (CCA) of various stages within PERL.
Model Code Code-FS Code-FS-$S$ Code-FS-$S$-VD
From C++ 86.52 86.31 (-0.21) 86.63 (+0.32) 87.22 (+0.59)
To C++ 90.97 91.08 (+0.11) 92.09 (+1.01) 91.97 (-0.12)
From Python 84.65 85.77 (+1.12) 85.38 (-0.38) 85.89 (+0.51)
To Python 85.62 85.24 (-0.38) 86.11 (+0.87) 86.59 (+0.48)
From Java 89.33 88.42 (-0.91) 89.71 (+1.29) 89.95 (+0.24)
To Java 83.91 84.48 (+0.57) 83.53 (-0.95) 84.50 (+0.97)
Average 86.83 86.88 (+0.05) 87.24 (+0.36) 87.69 (+0.45)
Table 8. Test Computational Accuracy (TCA) of various stages within PERL.
Model Code Code-FS Code-FS-$S$ Code-FS-$S$-VD
From C++ 82.37 82.37 (+0.00) 82.38 (+0.01) 83.23 (+0.84)
To C++ 85.01 85.87 (+0.86) 87.05 (+1.18) 86.62 (-0.43)
From Python 78.80 80.50 (+1.70) 80.31 (-0.19) 80.72 (+0.41)
To Python 79.31 79.42 (+0.11) 80.17 (+0.75) 80.93 (+0.76)
From Java 84.09 83.87 (-0.21) 84.95 (+1.07) 85.16 (+0.22)
To Java 80.94 81.46 (+0.52) 80.42 (-1.04) 81.56 (+1.15)
Average 81.75 82.25 (+0.50) 82.55 (+0.30) 83.03 (+0.49)
5. Conclusion
Appendix A Construction of Executability Representation
@@
Figure 5. Example C++ code used for encoding executability representations.
A.1. Variable-dependency Representation
@@
Figure 6. Data flow graph of the example code.
@@
Figure 7. Variable-dependency representation of the example code.
A.2. Syntactic-structure representation
@@
Figure 8. Abstract syntax tree of the example code.
@@
Figure 9. Syntactic-structure representation of the example code.
Appendix B Construction of XLCoST-Instruct
@@
ExeCoder selected parallel data from three programming languages: C++, Python, and Java, to construct a fine-tuning dataset, as these are the most widely used programming languages, and their translation tasks can effectively assess the code translation capabilities of LLMs. Here, the source code refers to complete programs rather than mere code snippets. Subsequently, we encoded three types of executability representations from the source code of XLCoST. For functional semantic representation, we encoded the natural language from the natural language-code pairs provided by XLCoST. For syntactic structure representation and variable dependency representation, we followed the methods outlined in Appendix A to encode from the source code. Based on this, a single data instance resulted in four different code representations. Since XLCoST is not specifically designed for instruction fine-tuning of LLMs, ExeCoder manually constructs instructions tailored for translation tasks, with more detailed information provided in Appendix C.
@@
Appendix C Example in XLCoST-Instruct

### v3

@@ 4 Experimental Evaluation
Table 1: Execution-based evaluation of ExeCoder and baseline models on TransCoder-test-X.
4.1 Experimental Setup
Dataset. We conducted comprehensive experiments on the widely used code translation public dataset TransCoder-test Roziere et al. (2020) to demonstrate the effectiveness of the ExeCoder. TransCoder-test contains solutions to programming problems collected from GeeksForGeeks, with each problem implemented in C++, Java, and Python, totaling 948 parallel samples. TransCoder-test is equipped with predefined unit test templates for over half of the translation pairs to evaluate whether the generated functions return the same output as the reference functions given the same input. Each unit test consists of ten test cases.
However, these unit test templates establish fixed parameter passing methods or return types that can only evaluate specific implementations. When the generated functions do not conform to the predefined implementations, the unit tests cannot execute as expected, even if the generated functions are functionally equivalent to the reference functions, resulting in unreliable evaluation. To address this, we enhanced TransCoder-test to ensure that the unit test results are capable of assessing the code translation ability. The enhanced test set is referred to as TransCoder-test-X, with detailed processing cases provided in Appendix G.
Match-based Metrics. The matching-based metrics aim to evaluate the quality of generated code through static code analysis. Following existing work Huang et al. (2023), we employed three metrics: Exact Match (EM), BLEU Papineni et al. (2002), and CodeBLEU Ren et al. (2020) for evaluation. These metrics can measure the n-gram overlap, syntactic structure, and semantic equivalence of the generated code.
Execution-based Metrics. The execution-based metrics aim to assess the quality of generated code through the execution status of the code. Following existing work Roziere et al. (2020), we utilize test computational accuracy (TCA) as a metric, which measures the proportion of code that passes unit tests. Furthermore, to further reflect the fine-grained execution status of the code, we have designed two additional metrics: Compilation Accuracy (CA) and Case Computational Accuracy (CCA), the former measuring the proportion of code that successfully compiles, and the latter measuring the proportion of code that passes case tests.
Baselines. To assess the effectiveness of ExeCoder, we compare it with state-of-the-art LLMs. Specifically, we use three advanced closed-source models as baselines: OpenAI’s GPT-3.5, GPT-4, and GPT-4-o Achiam et al. (2023). Additionally, we select five leading open-source code LLMs as baselines: Deepseek-Coder-6.7b-instruct (our base model) Guo et al. (2024), CodeLLama-7B Roziere et al. (2023), Magicoder-S-DS-6.7B Wei et al. (2024), Qwen2.5-Coder-7B Hui et al. (2024), and Wavecoder-ultra-6.7b Yu et al. (2024).
4.2 Experimental Results
Effectiveness of ExeCoder. To assess the effectiveness of ExeCoder in code translation task, we evaluated ExeCoder on the TransCoder-test-X, which includes translations among the three most commonly used programming languages: C++, Python, and Java. Tables 1 and 9 present the comparison results with baseline models.
Figure 4: Match-based evaluation of different variants of ExeCoder on TransCoder-test-X.
  • Code: Fine-tuning using only the source code.
Tables 2, 3, 4 and Figure 4 present our results, which indicate that the learning from each executability representation contributes to varying degrees of improvement. For execution-based metrics, the inclusion of the three different types of executability representations resulted in an average increase of 0.41%, 1.64%, and 1.21%. For match-based metrics, the integration of the three different types of executability representations led to an average increase of 0.39%, 5.56%, and 5.24%. These advantages indicate that executability representations can provide execution status beyond code semantics, enhancing the cross-language understanding capability of LLMs.
Model Code with FS with SS with VD
From C++ 91.13 91.56 (+0.43) 91.66 (+0.53) 91.14 (+0.01)
To C++ 96.36 96.47 (+0.11) 96.90 (+0.54) 97.11 (+0.75)
From Python 89.21 89.52 (+0.31) 91.20 (+1.99) 90.78 (+1.57)
To Python 90.41 91.59 (+1.18) 93.00 (+2.59) 92.89 (+2.48)
From Java 92.68 93.54 (+0.86) 95.05 (+2.37) 95.27 (+2.59)
To Java 86.25 86.56 (+0.31) 88.02 (+1.77) 87.19 (+0.94)
Average 91.01 91.54 (+0.53) 92.64 (+1.63) 92.39 (+1.38)
Table 2: Compilation Accuracy (CA) of different variants of ExeCoder.
Model Code with FS with SS with VD
From C++ 86.04 86.48 (+0.44) 87.33 (+1.29) 86.56 (+0.52)
To C++ 91.21 90.89 (-0.32) 91.68 (+0.47) 91.78 (+0.57)
From Python 84.27 84.28 (+0.01) 85.37 (+1.10) 84.79 (+0.52)
To Python 83.11 83.96 (+0.85) 85.83 (+2.72) 85.59 (+2.48)
From Java 87.13 87.43 (+0.30) 89.38 (+2.25) 89.38 (+2.25)
To Java 83.11 83.33 (+0.22) 84.57 (+1.46) 83.36 (+0.25)
Average 85.81 86.06 (+0.25) 87.36 (+1.55) 86.91 (+1.10)
Table 3: Case Computational Accuracy (CCA) of different variants of ExeCoder.
Model Code with FS with SS with VD
From C++ 81.62 82.17 (+0.55) 82.90 (+1.28) 82.16 (+0.54)
To C++ 85.22 85.12 (-0.10) 86.19 (+0.97) 86.08 (+0.86)
From Python 78.28 78.60 (+0.32) 80.08 (+1.80) 78.93 (+0.65)
To Python 76.40 77.59 (+1.19) 79.31 (+2.91) 79.20 (+2.80)
From Java 81.82 82.25 (+0.43) 83.98 (+2.16) 84.09 (+2.27)
To Java 80.10 80.31 (+0.21) 81.46 (+1.36) 79.90 (-0.20)
Average 80.58 81.01 (+0.43) 82.32 (+1.74) 81.73 (+1.15)
Table 4: Test Computational Accuracy (TCA) of different variants of ExeCoder.
Effectiveness of Progressive Executability Representation Learning. We conduct an ablation study to evaluate the effectiveness of the PERL strategy. Specifically, we assess the impact of each fine-tuning phase in PERL on the model’s code translation capability. Tables 5, 6, and 7 present our results. We also conducted a more detailed ablation study on PERL, with details provided in Appendix A due to space limitations.
Overall, ExeCoder’s meticulously designed PERL strategy significantly enhances the code translation performance through the learning of executability representations. Compared to models fine-tuned solely on source code, the inclusion of executability representations raises the upper limits of model capability, with three execution-based metrics improving by 0.53%, 0.85%, and 1.28%, respectively. Furthermore, the learning of specific executability representations at each stage improves the model’s cross-language understanding. Relative to the previous stage, each fine-tuning stage further improved the key execution-based metric TCA by 0.50%, 0.30%, and 0.49%, respectively.
Model Code Code-FS Code-FS-$S$ Code-FS-$S$-VD
From C++ 91.55 91.13 (-0.42) 91.15 (+0.02) 91.56 (+0.41)
To C++ 96.36 96.36 (+0.00) 97.11 (+0.75) 97.00 (-0.11)
From Python 89.73 90.78 (+1.05) 90.58 (-0.20) 91.00 (+0.42)
To Python 93.00 92.56 (-0.44) 93.53 (+0.97) 93.53 (+0.00)
From Java 95.16 94.62 (-0.54) 95.37 (+0.75) 95.48 (+0.11)
To Java 87.08 87.60 (+0.52) 86.46 (-1.14) 87.50 (+1.04)
Average 92.15 92.18 (+0.03) 92.37 (+0.19) 92.68 (+0.31)
Table 5: Compilation Accuracy (CA) of various stages within PERL.
Model Code Code-FS Code-FS-$S$ Code-FS-$S$-VD
From C++ 86.52 86.31 (-0.21) 86.63 (+0.32) 87.22 (+0.59)
To C++ 90.97 91.08 (+0.11) 92.09 (+1.01) 91.97 (-0.12)
From Python 84.65 85.77 (+1.12) 85.38 (-0.38) 85.89 (+0.51)
To Python 85.62 85.24 (-0.38) 86.11 (+0.87) 86.59 (+0.48)
From Java 89.33 88.42 (-0.91) 89.71 (+1.29) 89.95 (+0.24)
To Java 83.91 84.48 (+0.57) 83.53 (-0.95) 84.50 (+0.97)
Average 86.83 86.88 (+0.05) 87.24 (+0.36) 87.69 (+0.45)
Table 6: Case Computational Accuracy (CCA) of various stages within PERL.
Model Code Code-FS Code-FS-$S$ Code-FS-$S$-VD
From C++ 82.37 82.37 (+0.00) 82.38 (+0.01) 83.23 (+0.84)
To C++ 85.01 85.87 (+0.86) 87.05 (+1.18) 86.62 (-0.43)
From Python 78.80 80.50 (+1.70) 80.31 (-0.19) 80.72 (+0.41)
To Python 79.31 79.42 (+0.11) 80.17 (+0.75) 80.93 (+0.76)
From Java 84.09 83.87 (-0.21) 84.95 (+1.07) 85.16 (+0.22)
To Java 80.94 81.46 (+0.52) 80.42 (-1.04) 81.56 (+1.15)
Average 81.75 82.25 (+0.50) 82.55 (+0.30) 83.03 (+0.49)
Table 7: Test Computational Accuracy (TCA) of various stages within PERL.
5 Conclusion
Limitations
A primary limitation of this work is identified. ExeCoder employs static analysis tools to derive executability representations from source code, and these representations often possess extensive contextual information. Although the number of tokens during the fine-tuning process stays within the model’s context window, it nonetheless elevates the relevant fine-tuning expenditures, consequently diminishing ExeCoder’s scalability to a certain extent. Addressing the reduction of these costs falls outside the purview of the current study and is not the central topic of this paper.
Appendix A Ablation of PERL
we conducted further ablation studies on PERL to demonstrate its necessity and effectiveness. Specifically, we adopted the single-stage mixed training approach and reversed the order of the various representations in PERL, with the evaluation results presented in the table 8. The results show that both training with mixed representations and altering the order of the various representations had varying degrees of impact on the performance of ExeCoder. The training method PERL, which aligns with programming learning theory, achieved optimal performance.
Compilation Accuracy (CA)
Model NL-AST-DFG (ExeCoder) NL-DFG-AST AST-NL-DFG AST-DFG-NL DFG-AST-NL DFG-NL-AST Shuffle
From C++ 91.559 91.354 91.347 91.243 91.782 91.020 91.861
To C++ 97.002 96.681 96.681 96.574 96.467 95.824 96.253
From Python 90.995 90.579 89.948 89.844 89.948 89.410 90.460
To Python 93.534 93.534 93.427 93.427 93.103 92.888 93.211
From Java 95.482 95.053 95.375 95.268 94.299 95.053 95.268
To Java 87.500 86.771 86.563 86.354 86.458 86.771 88.125
Average 92.679 92.329 92.223 92.118 92.010 91.828 92.530
Case Computational Accuracy (CCA)
Model NL-AST-DFG (ExeCoder) NL-DFG-AST AST-NL-DFG AST-DFG-NL DFG-AST-NL DFG-NL-AST Shuffle
From C++ 87.221 87.145 86.892 86.548 86.880 86.247 87.616
To C++ 91.970 91.842 91.713 91.842 91.028 90.664 90.867
From Python 85.891 85.542 84.940 84.807 84.335 84.334 84.682
To Python 86.590 86.396 86.353 85.868 85.017 85.491 85.847
From Java 89.948 89.186 89.745 89.615 88.184 89.251 89.218
To Java 84.500 83.635 83.510 83.260 83.354 83.677 84.802
Average 87.687 87.291 87.192 86.990 86.466 86.611 87.172
Test Computational Accuracy (TCA)
Model NL-AST-DFG (ExeCoder) NL-DFG-AST AST-NL-DFG AST-DFG-NL DFG-AST-NL DFG-NL-AST Shuffle
From C++ 83.226 82.813 83.226 82.583 82.475 82.148 83.308
To C++ 86.617 86.938 86.617 86.831 85.760 85.760 85.439
From Python 80.717 80.630 79.991 79.678 78.935 79.557 79.340
To Python 80.927 80.172 80.603 79.849 78.556 79.095 79.310
From Java 85.163 83.980 84.733 84.732 83.011 84.088 83.872
To Java 81.563 80.313 80.729 80.313 80.104 80.938 81.771
Average 83.035 82.474 82.650 82.331 81.473 81.931 82.173
Table 8: Additional Ablation of PERL.
Appendix B Match-based evaluation
@@
Table 9: Match-based evaluation of ExeCoder and baseline models on TransCoder-test-X.
Appendix C Implementation Details
We present detailed implementation details. Deepseek-Coder-6.7b-instruct serves as our basic foundation model. To fine-tune the base model, we employed specific configurations, including a batch size of 128, a sequence length of 3076, a learning rate of 2e-6, a cosine learning rate scheduler, and fp16 mixed precision.
Appendix D Construction of Executability Representation
@@
Figure 5: Example C++ code used for encoding executability representations.
D.1 Variable-dependency Representation
@@
Figure 6: Data flow graph of the example code.
@@
Figure 7: Variable-dependency representation of the example code.
D.2 Syntactic-structure representation
@@
Figure 8: Abstract syntax tree of the example code.
@@
Figure 9: Syntactic-structure representation of the example code.
Appendix E Construction of XLCoST-Instruct
@@
ExeCoder selected parallel data from three programming languages: C++, Python, and Java, to construct a fine-tuning dataset, as these are the most widely used programming languages, and their translation tasks can effectively assess the code translation capabilities of LLMs. Here, the source code refers to complete programs rather than mere code snippets. Subsequently, we encoded three types of executability representations from the source code of XLCoST. For functional semantic representation, we encoded the natural language from the natural language-code pairs provided by XLCoST. For syntactic structure representation and variable dependency representation, we followed the methods outlined in Appendix D to encode from the source code. Based on this, a single data instance resulted in four different code representations. Since XLCoST is not specifically designed for instruction fine-tuning of LLMs, ExeCoder manually constructs instructions tailored for translation tasks, with more detailed information provided in Appendix F.
@@
Appendix F Example in XLCoST-Instruct